OpenStudy (anonymous):

What is the sum of the arithmetic sequence 174, 168, 162 …, if there are 31 terms? 2,604 2,635 2,666 2,697

OpenStudy (anonymous):

174 + [174 - 6(1)] + [174 - 6(2)] + [174 - 6(3)] + . . . + [174 - 6(30)] = (174)(31) - 6(1 + 2 + 3 + . . . + 30) = (174)(31) - (6)(30)(31)/2 That last part is derived from: 1 + 2 + 3 + . . . + n = n(n + 1)/2

OpenStudy (anonymous):

So, all you have to do is simplify: (174)(31) - (6)(30)(31)/2

OpenStudy (anonymous):

i.....ohhh

OpenStudy (anonymous):

thanks abunch bro

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @XAY_____

OpenStudy (anonymous):

thnx again

OpenStudy (anonymous):

Anytime!